# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v2

@@ 5.5 Optimal control of housekeeping EPR
Using the housekeeping EPR in eq.˜2, we formulate the optimal control problem for $\dot{\Sigma}_{hk}$ using the external control of $F_{\alpha}$ from an initial to a final value $F_{\gamma}^{i}$ to $F_{\gamma}^{f}$. Physically, this implies that the control parameters $\{\lambda_{E}\}$ of $E$ are fixed and the state-space distribution is in steady state. This leads to vanishing driving work and excess EPR contributions, respectively. The driving Lagrangian for the EPR of housekeeping then reads $\mathcal{L}_{drv}^{*}=\frac{1}{2}\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}\left(\dot{F}_{{\alpha}}\right)^{2}$. The mass for the driving reads, $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=T_{\alpha}^{\perp}\cosh{\left(F_{\alpha}/2\right)}+T_{\alpha}^{\perp}F_{\alpha}\sinh{\left(F_{\alpha}/2\right)}/4$. It is rewritten as $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=\frac{1}{2}\left(T_{\alpha}(A_{\alpha})+T_{\alpha}(A_{\alpha}^{\dagger})\right)+\frac{1}{4}\dot{\Sigma}_{\alpha}^{hk}$ with the total affinity $A_{\alpha}=F_{\alpha}+A_{\alpha}^{ex}$ and $A_{\alpha}^{\dagger}=F_{\alpha}-A_{\alpha}^{ex}$ being the affinity obtained by the time-reversal of the boundary term. Using the orthogonal symmetry of the non-equilibrium fluctuations, $T_{\alpha}(A_{\alpha})=T_{\alpha}(A_{\alpha}^{\dagger})$, the mass is reduced to, $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=T_{\alpha}+\dot{\Sigma}_{\alpha}^{hk}/4$. Hence, for driving the fEQ $F_{\alpha}$, the mass is proportional to the traffic $T_{\alpha}$ and the housekeeping EPR $\dot{\Sigma}_{\alpha}^{hk}$ due to the bidirectional transition $\alpha$, and increases exponentially as the system is further from equilibrium due to the hyperbolic scalings of $T_{\alpha}$ and $\dot{\Sigma}_{\alpha}$. This reveals that the higher housekeeping EPR and non-equilibrium fluctuations generate higher driving resistance. We use an exponential approximation of the hyperbolic functions, which is valid for fEQ systems, and use a previously computed closed-form analytical expression for the fEQ geodesic $\mathcal{G}^{fEQ}$ in eq.˜9.
@@
$\begin{split}\Sigma_{\tau}^{*hk}&=\frac{1}{2(2+\tau)}\left(\mathcal{G}^{fEQ}(F_{\alpha}^{f})-\mathcal{G}^{fEQ}(F_{\alpha}^{i})\right)^{2},\\ \mathcal{G}^{fEQ}(F_{\alpha})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{i})\\ &\hskip 28.45274pt+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{f}).\end{split}$ (40)
@@ 5.6 Multi-parameter optimal control of
We show the applicability of the multi-parameter optimal control. We incorporate the symmetric part of the transition, namely mobility $D_{\alpha}$, which was previously fixed. This implies $\{{\lambda}_{\alpha}^{i}\}=\{A_{\alpha},D_{\alpha}\},\forall\alpha\in\{\gamma^{\rightleftharpoons}\}$. Henceforth, we use the shorthand notation $a=A_{\alpha}$ and $d=D_{\alpha}$ for brevity. Using $\mathcal{L}^{*}$, the exact metric tensor for the simultaneous control of mobility $D_{\alpha}$ and affinity $A_{\alpha}$ reads
@@
Using eqs.˜41 and 26, the exact expression for the Christoffel symbols reads,
@@
The exact geodesic eq.˜25 for the optimal control of transition affinity and mobility is reduced to,
$\ddot{A}_{\alpha}+\Gamma_{aa}^{a}\>\dot{A}_{\alpha}^{2}=0$ (43a)
$\ddot{D}_{\alpha}+\Gamma_{aa}^{d}\>\dot{A}_{\alpha}^{2}+2\Gamma_{ad}^{d}\>\dot{A}_{\alpha}\dot{D}_{\alpha}=0$ (43b)
Equation˜43a for the optimal driving of the affinity is the same as eq.˜6, and does not depend on $D_{\alpha}$. Therefore, the optimal driving dynamics of the mobility $D_{\alpha}$ given by eq.˜43b are enslaved by the geodesic for $A_{\alpha}$. This implies that an optimal control problem for transition affinity and mobility can be reduced to two steps. First step, to solve the optimal control problem $A_{\alpha}$, as formulated in sections˜2 and 3, and obtain the geodesic $\mathcal{G}(A_{\alpha})$. Second step, plugging $\mathcal{G}(A_{\alpha})$ into eq.˜43b and solving ODE eq.˜43b for optimal driving of $D_{\alpha}$, and obtain the geodesic $\mathcal{G}(A_{\alpha},D_{\alpha})$. We cannot find an exact analytical solution for the geodesic $\mathcal{G}(A_{\alpha},D_{\alpha})$ due to the sophistication associated with solving coupled ODEs. However, the numerical computation of $\mathcal{G}(A_{\alpha},D_{\alpha})$ is always feasible.
@@ 6 Conclusion and Outlook
We consider a finite-time optimal control problem for driving the control parameters of discrete-state systems from the initial to the final value, such that the EPR for driving is minimised. Building upon the ‘minimum action principle’ (MinAP), we propose a framework for the generalized finite-time optimal control (GFTOC). The ‘minimum action principle’ allows for a variational formulation of the finite-time optimal control problem. To solve this problem, we develope a two-step solution. In the first step, assuming slow driving, we solve the slow driving optimal control problem by exploiting the Riemannian geometry induced in the manifold of control parameters. This approach unifies and generalizes the thermodynamic geometry and the OMTT for fEQ systems. Then, the dissipation-minimizing optimal paths are obtained by the geodesic, the minimum-distance path on the Riemannian manifold.
@@
Our framework opens up a plethora of practical applications in biology, chemistry, and nanoscale / mesoscale devices, where stochastic thermodynamics has been an experimentally tested theoretical paradigm [4]. Our framework opens the door for the experimental design, optimization, and control of such systems. The experimental verification of our framework awaits exploration [30]. The GFTOC framework is rather trivially extended to quantum systems, as the development of optimal control and speed limits for quantum systems is attributed to the underlying geometric structure and not to the quantum nature [77, 78, 79].

### v3

@@ 5.5 Optimal control of housekeeping EPR
Using the housekeeping EPR in eq.˜2, we formulate a novel and up-to-unresolved optimal control problem for $\dot{\Sigma}_{hk}$, using the external control of $F_{\alpha}$ from an initial to a final value $F_{\gamma}^{i}$ to $F_{\gamma}^{f}$. Physically, this implies that the control parameters $\{\lambda_{E}\}$ of $E$ are fixed and the state-space distribution is in steady state. This leads to vanishing driving work and excess EPR contributions, respectively. The driving Lagrangian for the EPR of housekeeping then reads $\mathcal{L}_{drv}^{*}=\frac{1}{2}\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}\left(\dot{F}_{{\alpha}}\right)^{2}$. The mass for the driving reads, $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=T_{\alpha}^{\perp}\cosh{\left(F_{\alpha}/2\right)}+T_{\alpha}^{\perp}F_{\alpha}\sinh{\left(F_{\alpha}/2\right)}/4$. It is rewritten as $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=\frac{1}{2}\left(T_{\alpha}(A_{\alpha})+T_{\alpha}(A_{\alpha}^{\dagger})\right)+\frac{1}{4}\dot{\Sigma}_{\alpha}^{hk}$ with the total affinity $A_{\alpha}=F_{\alpha}+A_{\alpha}^{ex}$ and $A_{\alpha}^{\dagger}=F_{\alpha}-A_{\alpha}^{ex}$ being the affinity obtained by the time-reversal of the boundary term. Using the orthogonal symmetry of the non-equilibrium fluctuations, $T_{\alpha}(A_{\alpha})=T_{\alpha}(A_{\alpha}^{\dagger})$, the mass is reduced to, $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=T_{\alpha}+\dot{\Sigma}_{\alpha}^{hk}/4$. Hence, for the fEQ driving of $F_{\alpha}$, the mass is proportional to the traffic $T_{\alpha}$ and the housekeeping EPR $\dot{\Sigma}_{\alpha}^{hk}$ due to the bidirectional transition $\alpha$, and increases exponentially as the system is further from equilibrium due to the hyperbolic scalings of $T_{\alpha}$ and $\dot{\Sigma}_{\alpha}$. This reveals that the higher housekeeping EPR and non-equilibrium fluctuations generate higher driving resistance. Therefore, we use an exponential approximation of the hyperbolic functions, which is valid for fEQ systems, and use a previously computed closed-form analytical expression for the fEQ geodesic $\mathcal{G}^{fEQ}$ in eq.˜10.
@@
$\begin{split}\Sigma_{\tau}^{*hk}&=\frac{1}{2(2+\tau)}\left(\mathcal{G}^{fEQ}(F_{\alpha}^{f})-\mathcal{G}^{fEQ}(F_{\alpha}^{i})\right)^{2},\\ \mathcal{G}_{\tau}^{fEQ}(F_{\alpha})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{i})\\ &\hskip 28.45274pt+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{f}).\end{split}$ (45)
@@ 5.6 Multi-parameter optimal control of
We show the applicability of the multi-parameter optimal control. We incorporate the symmetric part of the transition, namely mobility $D_{\alpha}$, which was previously fixed. This implies $\{{\lambda}_{\alpha}^{i}\}=\{A_{\alpha},D_{\alpha}\},\forall\alpha\in\{\gamma^{\rightleftharpoons}\}$. Henceforth, we use the shorthand notation $a=A_{\alpha}$ and $d=D_{\alpha}$ for brevity. Using $\mathcal{L}^{*}$, we pose a novel finite-time optimal control problem for the simultaneous control of mobility $D_{\alpha}$ and affinity $A_{\alpha}$, Then, the exact metric tensor reads
@@
Using eqs.˜46 and 31, the exact expression for the Christoffel symbols reads,
@@
The exact geodesic eq.˜30 for the optimal control of transition affinity and mobility is reduced to,
$\ddot{A}_{\alpha}+\Gamma_{aa}^{a}\>\dot{A}_{\alpha}^{2}=0,$ (48a)
$\ddot{D}_{\alpha}+\Gamma_{aa}^{d}\>\dot{A}_{\alpha}^{2}+2\Gamma_{ad}^{d}\>\dot{A}_{\alpha}\dot{D}_{\alpha}=0.$ (48b)
Equation˜48a for the optimal driving of the affinity is the same as eq.˜7, and does not depend on $D_{\alpha}$. Therefore, the optimal driving dynamics of the mobility $D_{\alpha}$ given by eq.˜48b are enslaved by the geodesic for $A_{\alpha}$. This implies that an optimal control problem for transition affinity and mobility can be reduced to two steps. First step, to solve the optimal control problem $A_{\alpha}$, as formulated in sections˜2 and 3, and obtain the geodesic $\mathcal{G}(A_{\alpha})$. Second step, plugging $\mathcal{G}(A_{\alpha})$ into eq.˜48b and solving ODE eq.˜48b for optimal driving of $D_{\alpha}$, and obtain the geodesic $\mathcal{G}(A_{\alpha},D_{\alpha})$. We cannot find an exact analytical solution for the geodesic $\mathcal{G}(A_{\alpha},D_{\alpha})$ due to the sophistication associated with solving coupled ODEs. However, the numerical computation of $\mathcal{G}(A_{\alpha},D_{\alpha})$ is always feasible.
@@ 6 Conclusion and Outlook
We consider a finite-time optimal control problem for driving the control parameters of discrete-state systems from the initial to the final value, such that the EPR for driving is minimised. Building upon the ‘minimum action principle’ (MinAP), we propose a framework for the generalized finite-time optimal control (GFTOC). The ‘minimum action principle’ allows for a variational formulation of the finite-time optimal control problem. To solve this problem, we develop a two-step solution. In the first step, assuming slow driving, we solve the slow driving optimal control problem by exploiting the Riemannian geometry induced in the manifold of control parameters. This approach unifies and generalizes the thermodynamic geometry and the Optimal TT for fEQ systems. Then, the dissipation-minimizing optimal paths are obtained by the geodesic, the minimum-distance path on the Riemannian manifold.
@@
Our framework opens up a plethora of practical applications in biology, chemistry, population dynamics, and nanoscale / mesoscale devices, where stochastic thermodynamics has been an experimentally tested theoretical paradigm [4]. Our framework opens the door for the experimental design, optimization, and control of such systems. The experimental verification of our framework awaits exploration [37]. The GFTOC framework is rather trivially extended to quantum systems, as the development of optimal control and geometric speed limits for quantum systems is attributed to the underlying geometric structure and not to the quantum nature [91, 92, 93]. More broadly, the GFTOC methodology developed here in stochastic thermodynamics applies to other interdisciplinary domains that rely on Riemannian geometry as their mathematical foundation, including information geometry [38], in which the role of ‘kinks’ warrants further investigation.
@@ Acknowledgements.
Note added. —  After the first publicly available draft of this work was released, we noted that the GFTOC framework solves several major open problems identified in the ‘Perspectives and Conclusions’ section of the recent review Ref.[12].